40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1007     Entry time: Mon Sep 29 15:09:36 2008
Author: steve 
Type: Update 
Category: PSL 
Subject: almost 4 yrs plot of power & temps 
The water chiller is normally running 1.5 C warmer than the laser head temp.
When control room temp is stable and PEM-count_temp is stable we can expect the head temp to be stable 20.0 C

PSL-126MOPA_HTEMP is running warmer in the last ~40 days

The ifo arm thermostate temp settings were raised by 2 F on 8-11-08
Attachment 1: 3.5y.jpg  117 kB  | Hide | Hide all
3.5y.jpg
ELOG V3.1.3-